At a certain instant the dimensions of the rectangle are 8 and 12 feet, and they are increasing at the rates 3 and 2 feet per second, respectively. How fast is the area changing?


1
Expert's answer
2021-11-15T20:07:15-0500

"x(t) =" the length of the rectangle at time t

"y(t) =" the width of the rectangle at time t

The area of the rectangle at time t is:

"A = xy"

The rate at which the area of the rectangle changes is:

"dA / dt = ( dx / dt )( y ) + ( x )( dy / dt)\\

x = 8\\

y = 12\\

dx / dt = 3\\

dy / dt = 2\\

dA / dt = ( 3 )( 12 ) + ( 8 )( 2 ) = 36 + 16 = 52\ ft^{2 }/ s\\"